you are driving on a highway to a town that is 300 kilometers from your home.after 30 minutes you pass a free way exit that you know is 50 kilometers from your home.Assuming that you continue at the same constant speed, how long will it take for this entire trip?

Answer by stanbon(48568)

you are driving on a highway to a town that is 300 kilometers from your home.after 30 minutes you pass a free way exit that you know is 50 kilometers from your home.Assuming that you continue at the same constant speed, how long will it take for this entire trip?
----------
(300-50)km/30 minutes = 50/x minutes
x = 30*50/250
x = 6 minutes
---------
Cheers,
Stan H.